I'm in the 8th grade and I don't understand the following problem:
Determine all the values of x and y such that xy, x/y and x-y are all equal.
I know that x and y can not equal zero.
Thanks you for your help.

Answer by venugopalramana(3286)

Determine all the values of x and y such that xy, x/y and x-y are all equal.
LET US TAKE THE FIRST 2
XY=X/Y
Y^2=1
Y=1 OR -1
NOW LET US TRY THESE VALUE IN THE OTHER AND CHECK
XY=X-Y
CASE 1....Y=1
X=X-1....IMPOSSIBLE
CASE2...Y=-1
-X=X+1
-2X=1X=-1/2...SO THE POSSIBLE SOLUTION IS
X=-1/2 AND Y=-1...CHECK...WE GET
XY=-(1/2)*(-1)=1/2
X/Y=-(1/2)/(-1)=1/2
X-Y=-(1/2)-(-1)=1/2...OK...